# Improving Precision in Kinematic Weak Lensing with MIRoRS: Model-Independent Restoration of Reflection Symmetries

### v1

@@ 4.3 Illustris Data
Model Inclination $(^{\circ})$ Mean Median $\sigma$
Combined All -0.007 -0.003 0.039
$20$0.009 -0.003 0.047
$40$0.008 -0.003 0.040
$60$0.005 -0.003 0.027
Kinematic All -0.009 -0.006 0.050
$20$0.011 -0.008 0.060
$40$0.010 -0.005 0.049
$60$0.007 -0.005 0.039
Donet (2023) All -0.017 -0.010 0.078
$20$0.020 -0.012 0.088
$40$0.018 -0.010 0.080
$60$0.015 -0.009 0.063
Figure 9: Comparison between the combined model and the results from DW23. The nonparametric uncertainty of $0.039$ is a factor of two improvement over their parametric model.
@@
We analyze this Illustris data set and compare with the results from DW23 on the same data, shown in Figure 9. In that paper, they fit their distribution with a double Gaussian profile with a narrow “core”. Our method largely eliminates the broad wings seen in their distribution, such that our overall scatter is comparable to the scatter in the core of their distribution. While DW23 cites this low core uncertainty at $0.025$ to $0.036$, their total uncertainty is between $0.089$ and $0.062$, depending on inclination. Across the same inclinations, we report a total uncertainty of $0.027$ to $0.047$ with an overall uncertainty of $\sigma=0.039$ across the entire sample.
We can also see a slight bias towards negative shear, with all mean and median values being negative. The standard error in the mean of $0.001$ when including all inclinations. With mean and median values of $-0.003$ and $-0.007$, this result is significant at 3–7 sigma and matches the findings of DW23. As in that paper, we oriented all galaxies with a clockwise rotation (Z-wise orientation of spiral arms). This could produce the observed systematic bias. In a real observation averaged over an ensemble of galaxies of arbitrary orientation, we expect the bias to vanish. However, the bias could be considered in individual observations.
4.4 Moran’s I-test
Figure 10: Relationship between Moran’s I and p-value. All statistically significant p-values, $p\leq 0.001$, correspond to relatively large I-values.
@@ 4.4 Moran’s I-test
Finally, we use Moran’s I-test to filter the results based on coherent patterns in the residuals. The PySAL.esda package returns both an I-value and a p-value for the null hypothesis of complete spatial randomness. In Figure 10, we see a definite trend in the relationship between the p and I-values, with higher I-values corresponding to lower p-values.
@@
A good fit is one that exhibits spatial randomness, corresponding to an I-value near zero. Given the strong relationship between the p and I-values, we choose to filter the results based on p-value alone, rejecting results with $p\leq 0.001$. We can validate this choice by examining how the I and p-values relate to our fit for $\gamma_{\times}$, Figure 11. There is a definitive trend for large value p-values and small I-values when $\gamma_{\times}$ is near its true value of 00.
Table 4: Moran’s I filtered results.
Inclination $(^{\circ})$ Mean Median $\sigma$
All -0.007 -0.002 0.028
$20$0.009 -0.002 0.033
$40$0.007 -0.003 0.020
$60$0.003 -0.002 0.019
@@
Comparing the results of the Illustris data fit with the combined model and the same results filtered with the Moran’s I-test is given in Figure 12. The 358 halos at 3 inclinations each produce a total $1074$ mock observations. Of these, $472$ were filtered out, leaving $602$ with no consideration for inclination angle. The uncertainty in the sample was reduced from $0.039$ to $0.028$ with improvements at all inclination levels, shown in Table 4.
Figure 11: Results from Moran’s I-test with fitted values for $\gamma_{\times}$. Correlation between high values of Moran’s-I (low p-values) with large $\gamma_{\times}$ show spatial correlation within the residuals can indicate a poor fit.
Figure 12: Histogram of results after filtering to remove sources with a p-value in Moran’s I of $p\leq 0.001$. This filtering reduced the sample size from the original $1074$ ($358$ halos at three inclinations each) to $602$ and produced and improvement on the uncertainty to $\sigma=0.028$.
Figure 13: Visualization of Halo 491282. The data and model (left four panels) are observed at an inclination of $20^{\circ}$. The set of two panels on the right show the stellar and gas column densities of the galaxy observed face-on. The prominent substructure is unable to be modeled using symmetry operations and show as a pattern in the resulting residual after fitting with a shear parameter of $\gamma_{\times}=-0.0396$. This halo was filtered out with an I-value of $0.6435$. All images are 8 kpc across.
@@
In a real observation, the most likely cause of coherent patterns in the residual will be unmodeled physical effects. Features such as warps or bars will not display the same symmetries as an ideal disk. Filtering with Moran’s I-test can help to identify targets that are poorly fit due to these features. We choose to simply eliminate these candidates from our analysis, but it is conceivably possible to eliminate the problematic features through creative masking to select the most symmetric regions of the galaxy.
@@
We examine the halo from our sample with the largest Moran I-value. Halo $491282$ was fit with a shear parameter of $\gamma_{\times}=-0.0396$ and a Moran I-value of $0.6435$ and p-value of $0.001$ at an inclination of $20^{\circ}$. As can be seen in Figure 13, there are noticeable unmodeled features visible in the residual. While the kinematic and photometric data both look fairly innocuous, upon closer inspection of the stellar and gas column densities available from Illustris, there is distinct substructure present. These features are not symmetric and, thus, we are unable to precisely fit them with our methods.
5 Summary and Conclusion
@@
We first characterized our approach using idealized mock data, varying parameters for shear, position angle, inclination, and noise to explore the robustness and limitations of our method. Our nonparametric kinematic model showed improved precision over traditional parametric models, particularly when combined with photometric data. The addition of applied $\gamma_{+}$ increases scatter in the fit for $\gamma_{\times}$ with the combined model. However, for values of $\gamma_{+}$ up to 0.2, the uncertainty remains an improvement on both kinematic and nonparametric approaches tested without $\gamma_{+}$ applied.
Applying our model to a realistic dataset of 358 halos from the Illustris TNG simulations, we achieved a significant reduction in the uncertainty of $\gamma_{\times}$, demonstrating an overall uncertainty of $0.039$. These results represent a substantial improvement over a previous parametric approach by DW23, which reported uncertainties around 0.078 on the same dataset.
Additionally, we introduced a method for identifying and discarding outliers using Moran’s I-test for spatial autocorrelation. This technique effectively filtered out halos with coherent patterns in their residuals, reducing the overall uncertainty to $0.028$.
Our findings highlight the potential of integrating kinematic and photometric data in weak-lensing studies, offering a more precise measurement of weak-lensing shear.
6 Acknowledgements
We thank Jean Donet for his advice and assistance on the project, as well as his work in supplying all of the processed Illustris data files.

### v2

@@ IV.3 Illustris Data
Model Inclination (deg) Mean Median $\sigma$ $\sigma_{\mathrm{core}}$
Filtered All -0.007 -0.002 0.028 0.014
2020 -0.009 -0.002 0.033 0.012
$40$ -0.007 -0.003 0.030 0.013
$60$ -0.003 -0.002 0.019 0.012
Combined All -0.007 -0.003 0.039 0.016
2020 -0.009 -0.003 0.047 0.012
$40$ -0.008 -0.003 0.040 0.016
$60$ -0.005 -0.003 0.027 0.017
Kinematic All -0.009 -0.006 0.050 0.031
2020 -0.011 -0.008 0.060
$40$ -0.010 -0.005 0.049 0.033
$60$ -0.007 -0.005 0.039
Donet (2023) All -0.017 -0.010 0.078 0.030
2020 -0.018 -0.012 0.089 0.036
$40$ -0.015 -0.010 0.076 0.029
$60$ -0.014 -0.009 0.062 0.025
DW23 analyzed the same Illustris data set, allowing for a direct comparison between methods, with their results shown in blue in Figure 13. In that paper, they fit their distribution with a double Gaussian profile with a broad component plus a narrow core. Our method largely eliminates the broad wings seen in their distribution, such that our overall scatter is comparable to the scatter in the core of their distribution. While DW23 cites this low core uncertainty at 0.0250.0250.0360.036, their total uncertainty is between $0.089$ and $0.062$, depending on inclination. Across the same inclinations, we report a total uncertainty of 0.0270.0270.0470.047 with an overall uncertainty of $\sigma=0.039$ across the entire sample. Our comparable core uncertainties are roughly half the overall, at $0.016$ for the combined model across all inclinations (Table 3).
Comparing the two nonparametric models, shown in purple and green in Figure 13, we see the combined model shows a distinct improvement over the kinematic model. Total uncertainty in the fit of $\gamma_{\times}$ with the kinematic model is $0.05$ considering all inclination angles and improves to $0.039$ at an inclination of $60^{\circ}$. The combined model has an uncertainty of $0.039$, improving to $0.027$ at $i=60^{\circ}$.
@@
[…]
IV.4 Moran’s $\mathrm{I}$ Test
Figure 14: Moran’s II analysis of the Illustris sample. Top panel: fraction of realizations rejected as a function of fitted $\gamma_{\times}$ when applying a significance threshold of $p\leq 0.001$. Bottom panel: Moran’s II vs. fitted $\gamma_{\times}$ for all 1074 mock observations, with hexagonal binning colored by $\log p$–value. The red dashed line marks the rejection cutoff ($p\leq 0.001$), which removes 472 cases while retaining 602. Large values of Moran’s II (corresponding to low pp–values) correlate with large inferred $|\gamma_{\times}|$, indicating that spatially coherent residuals are associated with biased shear estimates.
@@ IV.4 Moran’s $\mathrm{I}$ Test
Finally, we use Moran’s $I$ test to filter the Illustris results based on coherent patterns in the residuals. The PySAL.esda package returns both an $I$-value and a $p$-value for the null hypothesis of complete spatial randomness. In Figure 14, we see a definite trend in the relationship between the $p$- and $I$-values, with higher $I$-values corresponding to lower $p$-values.
@@
A good fit is one that exhibits spatial randomness, corresponding to an I-value near zero. Given the strong relationship between the p and I-values, we choose to filter the results based on $p$-value alone, rejecting results with $p\leq 0.001$. We can validate this choice by examining how the I and $p$-values relate to our fit for $\gamma_{\times}$, Figure 14. There is a definitive trend for large value $p$-values and small I-values when $\gamma_{\times}$ is near its true value of 0.
@@
Filtering the results of the Illustris data fit with the combined model, shown in red in Figure 13, produces a noticeable improvement in the scatter. The 358 halos at three inclinations each produce a total $1074$ mock observations. Of these, $472$ were filtered out, leaving $602$ with no consideration for inclination angle. The uncertainty in the sample was reduced from $0.039$ to $0.028$ with improvements at all inclination levels, shown in Table 3.
The final distribution of fitted $\gamma_{\times}$ values, after filtering, is shown in red in Figure 13. Across all inclinations, the filtered core uncertainty is $0.014$, half the total uncertainty. The negative bias in the fit remains and dominates the fit of the broad component in all but the $i=60^{\circ}$ case. These galaxies can be identified in Figure 14 as the cluster of galaxies with $\gamma_{\times}\approx-0.1$ and $I<0.1$. It is also evident that these galaxies would remain in the filtered sample for any small adjustment in the filtering threshold.
@@
In a real observation, the most likely cause of coherent patterns in the residual will be unmodeled physical effects. Features such as warps or bars will not display the same symmetries as an ideal disk. Filtering with Moran’s $I$ test can help to identify targets that are poorly fit due to these features. We choose to simply eliminate these candidates from our analysis, but it is conceivably possible to eliminate the problematic features through creative masking to select the most symmetric regions of the galaxy.
Figure 15: Visualization of Halos 491282 and 483113 and comparison of $\gamma_{\times}$ fit with the MIRoRS model and with the methods of DW23. Top panel: results for Halo 491282. The four left panels show the stellar and gas column densities from Illustris along with the derived velocity and intensity fields. The set of four panels on the right shows the models and residuals from fitting with MIRoRS and a parametric algorithm based on DW23. The prominent substructure is unable to be modeled using symmetry operations and appears as a pattern in the resulting residual after fitting with a shear parameter of $\gamma_{\times}=-0.0396$ with MIRoRS and $\gamma_{\times}=-0.0989$ in DW23. This halo was filtered out of the MIRoRS sample with an I-value of $0.6435$ ($I=0.7004$ for the parametric model). Bottom panel: results for Halo 483113. The Moran test did not identify any correlated residuals in the MIRoRS fit with a value of $I=0.0601$. Shear was fit moderately well with MIRoRS ($\gamma_{\times}=-0.0268$), but the fit from DW23 ($\gamma_{\times}=-0.166$) was further from the expected value of zero. All images are 10 kpc across.
@@
We examine the halo from our sample with the largest Moran $I$-value. Halo $491282$, shown in the top panels of Figure 15, was fit with a shear parameter of $\gamma_{\times}=-0.0396$ with a Moran $I$-value of $0.6435$ and $p$-value of $0.001$ at an inclination of $20^{\circ}$. The right set of panels gives the model and residuals for both the MIRoRS method and the parametric approach implemented in DW23. There are noticeable unmodeled features visible in the residuals of both fits. While the kinematic and photometric data both look fairly innocuous, upon closer inspection of the stellar and gas column densities available from Illustris (far left panels), there is distinct substructure present. These features are not symmetric and, thus, we are unable to precisely fit them with our methods. The results from DW23 produced a fit further from the expected value at $\gamma_{\times}=-0.0989$ and display more pronounced residuals.
In comparison, we also examined a system that had no correlated residuals identified ($I=0.0601$) when fit with MIRoRS. Halo 483113 at $i=40^{\circ}$ was fit with a shear parameter of $\gamma_{\times}=-0.0268$ with MIRoRS while DW23 found $\gamma_{\times}=-0.166$. We also performed the Moran’s test with the residual produced by the parametric model and find the statistic larger, but not meeting our rejection threshold ($I=0.094$, $p=0.014$). Overall, this is an excellent example of the strengths of the MIRoRS method. Using our rejection threshold, both methods would be considered a successful fit, yet the shear parameter found with the parametric method is far from the expected truth.
V Summary and Conclusion
@@
We first characterized our approach using idealized mock data, varying parameters for shear, position angle, inclination, and noise to explore the robustness and limitations of our method. Our nonparametric kinematic model showed improved precision over traditional parametric models, particularly when combined with photometric data. The addition of applied $\gamma_{+}$ increases scatter in the fit for $\gamma_{\times}$ with the combined model. However, for values of $\gamma_{+}$ up to 0.2, the uncertainty remains an improvement on both kinematic and nonparametric approaches tested without $\gamma_{+}$ applied. Further, we demonstrate the necessity to align the galaxy major axis (kinematic, or unlensed) with the detector $x$-axis. We also explored the effect of asymmetries in the velocity field with mock data created with a linearly varying kinematic position angle and find a small systematic bias, but no increase in scatter.
Applying our model to a realistic dataset of 358 halos from the Illustris TNG simulations, we achieved a significant reduction in the uncertainty of $\gamma_{\times}$, demonstrating an overall uncertainty of $0.039$ and core uncertainty of $0.016$. These results represent a substantial improvement over a previous parametric approach by DW23, which reported an overall uncertainty of $0.078$ and core uncertainty of $0.030$ on the same dataset.
Additionally, we introduced a method for identifying and discarding outliers using Moran’s I-test for spatial autocorrelation. This technique effectively filtered out halos with coherent patterns in their residuals, reducing the overall uncertainty to $0.028$ (core $0.014$).
Our findings highlight the potential of integrating kinematic and photometric data in weak lensing studies, offering a more precise measurement of weak lensing shear.
Acknowledgments
We thank Jean Donet for his advice and assistance on the project, as well as his work in supplying all of the processed Illustris data files. We would also like to thank the anonymous referee for providing constructive feedback.